express the solution in exact form
ln(-x)+ ln 4=ln(3x-9)

Answer by stanbon(48535)

ln(-x)+ ln 4=ln(3x-9)
ln[-4x] = ln(3x-9)
Take the anti-ln of each side:
-4x=3x-9
7x=9
x=9/7
=====================
Cheers,
Stan H.